if u=(2,-3,4) & v=(-1,2,0)& w=(5,-1,2) find the angle b/n (u-2v) & (u+2w)


1
Expert's answer
2022-03-14T07:37:31-0400

"(u-2v) = (2,-3,4) - 2\cdot(-1,2,0) = (4,-7,4)"

"(u+2w) = (2,-3,4) + 2\cdot(5,-1,2) = (12, -5, 8)"

The angle "\alpha" between two vectors "\vec{a}" and "\vec{b}" can be obtained as "\cos\alpha = \dfrac{\vec{a}\cdot\vec{b}}{|\vec{a}||\vec{b}|}."

"\cos\alpha = \dfrac{4\cdot12+(-7)\cdot(-5)+4\cdot8}{\sqrt{4^2+(-7)^2+4^2}\sqrt{12^2+(-5)^2+8^2}},\\
\cos\alpha = \dfrac{115}{9\cdot\sqrt{233}}, \\
\alpha = 33.2^\circ."